the sum of three consecutive odd integers is the same as five times the greatest of these integers what are the integers?
Answer by Alan3354(69443)

the sum of three consecutive odd integers is the same as five times the greatest of these integers what are the integers?
x + x+2 + x+4 = 5*(x+4)
3x+6 = 5x+20
-2x = 14
x = -7
-----
-7, -5 and -3